In a certain chemical process, a lab technician supplies 251 J of heat to a system. At the same time, 64.0 J of work are done on the system by its surroundings. What is the increase in the internal energy of the system?
Answer #1

Since heat and work are both forms of energy. The first law of thermodynamics states that

dE=dQ +dW =(Heat absorbed) + (work done on the system)

the increase in the internal energy of the system = 251 + 64 =315 J